OpenStudy (anonymous):

Find the limit of rational function as x->infinity and as x->-infinity f(x)=(2x+3)/(5x+7)

OpenStudy (anonymous):

\[\Large \lim_{x \to \pm \infty} \frac{2x + 3}{5x + 7} = \lim_{x \to \pm \infty} \frac{2 + \frac{3}{x}}{5 + \frac{7}{x}}\] Now as x gets closer and closer to infinity, what do \(\Large \frac{3}{x}\) and \(\Large \frac{7}{x}\) tend to?

OpenStudy (anonymous):

do they tend to get closer to 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh so 2/5 is left

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

lol I don't know why I didn't see that

OpenStudy (anonymous):

You can also think if it another way, if x gets really large, the +3 and +7 in the denominator start to matter very little (2*100000000 + 3 for example, the + 3 is not really that significant), so you're only interested in 2x/5x as x gets really large or small which is easy to see that it's 2/5

OpenStudy (anonymous):

in the denominator and numerator*

OpenStudy (anonymous):

ok got it thanks @Meepi

OpenStudy (anonymous):

Check out www.aceyourcollegeclasses.com you can get homework and make money answering questions and uploading homework tutorials.